OpenStudy (anonymous):

In a large bag of Skittles, each of the 5 colors (red, orange, yellow, blue, green) occurs with the same probability. You reach in and select 1 candy. Find P(yellow or orange)

OpenStudy (anonymous):

I'm thinking 0.40 since there are 2/5

hartnn (hartnn):

it is correct :) yellow or orange = 1+1 total =5 P = 2/5

OpenStudy (anonymous):

Yep saw your message right when I bumped it xD Thanks again!

hartnn (hartnn):

welcome :)